What would be the pH of a solution prepared by dissolving 19.2 g NaOH (MM = 40.0 g/mol) in enough water to make 1.05 L of solution?


1
Expert's answer
2021-02-25T03:45:12-0500

mol of NaOH = mass/MW = 19.2/40 = 0.48 mol

V = 1.05L

M = mol/V = 0.48/1.05

M =0.45714 M

pOH = -log(0.45714) = 0.33995

pH = 14-0.33995

pH = 13.66005